OpenStudy (anonymous):

1/a+3-4/a-3=2a/a^2-9 a=?

OpenStudy (anonymous):

a=5/9

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ a+3\end{matrix}\right)-\left(\begin{matrix}4 \\ a-3\end{matrix}\right)=\left(\begin{matrix}2a \\ a ^{2}-9\end{matrix}\right)\]

OpenStudy (anonymous):

5/9 was incorrect

OpenStudy (anonymous):

No solutions exist for your equation. 5/9 was wrong because the format you used in your first post was a little bit messed up.

OpenStudy (anonymous):

yeah it was sorry

OpenStudy (anonymous):

Assuming that is 1/(a+3) and 4/(a-3) and (2a)/(a^2-9) I ended up with the same 5/9 as Kons.

OpenStudy (anonymous):

\[(1/(a+3))*(a-3)/(a-3)-(4/(a-3))*(a+3)/(a+3)=(2a)/(a ^{2}-9)\] \[(a-3)/(a ^{2}-9)-4(a+3)/(a^{2}-9)=2a/(a^{2}-9)\] \[(a-3-4a-12)/(a ^{2}-9)=(2a)/(a ^{2}-9)\] Times both sides by (a^2-9) \[-3a-15=2a\] \[-15=5a\] a=-3

OpenStudy (anonymous):

Well I take back what I said. I missed applying a minus sign and found it typing out the problem :p

OpenStudy (anonymous):

thank you!